If the position vector of one end of a chord through the focus of the parabola y^2 = 8x is 1/2i + 2j, find the position vector of the other end.



1
Expert's answer
2021-07-30T14:40:45-0400

Given parabola "y^{2}=8x" , the position vector of one end of an focal chord is "\dfrac{1}{2}\vec{i}+2\vec{j}."

We know that we let points on parabola as


"(at^{2},2at)=(\dfrac{1}{2},2)"

The value of "a" for given parabola is "a=2".

So comparing the both sides of given equation , we get


"at^{2}=\dfrac{1}{2}, 2at=2=>t=\dfrac{1}{2}"


Any other point on the parabola will be "t_{1}."

Suppose the coordinates of the extremity P of the focal chord are "(at^2, 2a)," the coordinates of the other extremity Q of the focal chord are "(at_1^2, 2at_1)," and the coordinates of the focus "F" are "(a, 0)."

Then, PF and FQ,have the same slopes


"\dfrac{2at-0}{at^2-a}=\dfrac{2at_1-0}{at_1^2-a}"

"tt_1^2-t=t_1t^2-t_1"

"tt_1(t_1-t)=t-t_1"

We get


"tt_{1}=-1"


So we get the value of "t_{1}=-2" .

Then other point on focal chord of parabola is


"(at_1^{2},2at_1)=(8,-8)"


So position vector at the other end will be


"8\vec{i}-8\vec{j}"